OpenStudy (mimi_x3):

The elevation angles of the tree top from Kevin and Lisa standing on the ground are respectively 40° and 30°. The bearings of Kevin and Lisa from the tree foot are respectively due East and SE. If the tree is 20m high, how far apart are Kevin and Lisa, correct to the nearest metre?

OpenStudy (anonymous):

631.891 meters

OpenStudy (mimi_x3):

Um, how did you get the answer ?

OpenStudy (anonymous):

distance of Kevin from tree x = 20/tan(40)=23.835 distance of Lisa from tree x = 20/tan(30)=34.641  distance between Kevin and Lisa = 34.641^2 - 23.835^2 =631.891 meters